OpenStudy (anonymous):

Given a soda can with a volume of 36 and a diameter of 4, what is the volume of a cone that fits perfectly inside the soda can?

OpenStudy (anonymous):

can is considered as a cylinder..so for cylinder volume=pi*r*h and for cone volume=pi*r*h/3 so simply put values in formula of cone and get result.

OpenStudy (anonymous):

Volume of inscribed cone is 1/3 volume of the cylinder. 12 is wat u get

OpenStudy (anonymous):

Wait.. what's 12? Sorry. Didn't follow.

OpenStudy (anonymous):

12 is the volume of a cone that fits perfectly inside the soda can 36/3 is equal to 12

OpenStudy (anonymous):

Ohhh okay I see now. Thank you!